OpenStudy (anonymous):

write the sum of the given polynomial in simplest form. (4x2 - 3x - 7) + (3x2 - 2x + 3)

hartnn (hartnn):

combine like terms like combine 4x^2 and 3x^2 = 7x^2 can u combine others?

OpenStudy (anonymous):

3x and 2x, -7 and 3 ?

hartnn (hartnn):

-3x-2x = -5x and -7+3 = -4 so overall 7x^2-5x-4 ok?

OpenStudy (anonymous):

yea i got it thank.